# Contractive kinetic Langevin samplers beyond global Lipschitz continuity

### v1

@@ 1 Introduction
$\mathrm{d}L_{t}$ $=-\nabla u\left(L_{t}\right)\mathrm{d}t+\sqrt{\frac{2}{\beta}}\mathrm{\leavevmode\nobreak\ d}B_{t},\quad t\geq 0$
@@
and exploiting the fact that this SDE admits the target measure $\exp(-\beta u)$ as its unique invariant measure. Inspired by this observation Langevin-based algorithms have become a popular choice for sampling from distributions in high dimensional spaces. There has been a lot of work in providing non-asymptotic results for Unadjusted Langevin algorithm (ULA) under various assumptions such as Lipschitz continuity of $\nabla u$ and convexity of $u$. Under the assumption of convexity and gradient Lispchitz continuity important results are obtained in Dalalyan (2017); Durmus and Moulines (2017, 2019); Sabanis and Zhang (2019); Barkhagen et al. (2021), while in the non-convex case, under convexity at infinity or dissipativity assumptions, one may consult Cheng et al. (2018a) Majka et al. (2020), Erdogdu et al. (2022) for ULA while for the Stochastic Gradient variant (SGLD) important works are Raginsky et al. (2017), Chau et al. (2021), Zhang et al. (2023b).
More recently, starting with the work of Vempala and Wibisono (2019) important estimates have been obtained under the assumption that the target measure $\pi$ satisfies an isoperimetric inequality and the gradient of $u$ satisfies a global Lipschitz continuity, (Mou et al. (2022), Balasubramanian et al. (2022)). The latter assumption has also been relaxed to a weakly smooth assumption (with the gradient still satisfying a linear growth property), see Nguyen et al. (2021) and Erdogdu and Hosseinzadeh (2021).
Recent advances on sampling has led to a connection with optimization (Vempala and Wibisono (2019),Chewi and Stromme (2024)) where overdamped Langevin dynamics can be seen as an analogous to gradient descent. It is therefore natural to explore momentum-like algorithms (nes (1983)) for possible improvements. An example of such a method is the discretization of the underdamped/kinetic Langevin SDE given as
@@
where $\gamma>0$ is the friction coefficient and $B_{t}$ is a $d$-dimensional Brownian motion. Similarly to the overdamped Langevin SDE, the discretization of this diffusion can be used as both an MCMC sampler and non-convex optimizer ( Gao et al. (2021)) since under appropriate conditions, the Markov process $\left\{\left(\tilde{\theta}_{t},\tilde{V}_{t}\right)\right\}_{t\geq 0}$ has a unique invariant measure given by
@@
Consequently, the marginal distribution of (2) in $\theta$ is precisely the target measure $\pi$. This means that sampling from (2) in the extended space and then keeping the samples in the $\theta$ space defines a valid sampler for the density $\pi$. The underdamped Langevin MCMC has shown improved convergence rates in the case of convex and gradient Lipschitz smooth potential $u$, e.g see Cheng et al. (2018b), Dalalyan and Riou-Durand (2020),Monmarché (2021)Altschuler and Chewi (2024). In the non-convex setting important work has been done in Akyildiz and Sabanis (2020), Gao et al. (2018), Chau and Rasonyi (2019),Ma et al. (2021),Zhang et al. (2023a), while there has been a lot of effort in analyzing the contractive behaviour of these algorithms (Monmarché (2021), Leimkuhler et al. (2024)).
@@ 2.1 Blanket assumptions
$\langle h(x)-h(y),x-y\rangle\geq 2m|x-y|^{2}\quad x,y\in\mathbb{R}^{d}.$
@@
$\mathbb{E}|\theta_{0}|^{2k}+\mathbb{E}|V_{0}|^{2k}<\infty\quad\forall k\in\mathbb{N}.$
@@ 2.2.1 Stochastic exponential scheme
In this section we propose a modified version of the stochastic exponential scheme, an algorithm which was first developed in Cheng et al. (2018b) and analysed in depth in Dalalyan and Riou-Durand (2020) and Gao et al. (2021) under the assumption of Lipschitz continuity for the gradient. The creation of this algorithm is motivated by the fact that (1), after applying Itô’s formula to the product $e^{\gamma t}V_{t}$ and by following standard calculations, can be rewritten as
@@
  • for any fixed $j$, the random vectors $\left(\left(\boldsymbol{\Xi}_{j}\right)_{1},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{1}\right),\left(\left(\boldsymbol{\Xi}_{j}\right)_{2},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{2}\right),\ldots,\left(\left(\boldsymbol{\Xi}_{j}\right)_{d},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{d}\right)$ are iid with the covariance matrix
@@
$\mathcal{L}(\bar{Y}^{\lambda}_{n},\bar{x}^{\lambda}_{n})=\mathcal{L}(\tilde{Y}_{n\lambda},\tilde{x}_{n\lambda}),\quad\forall n\in\mathbb{N}.$
@@ 3 Monotonicity preserving taming scheme
Taming schemes have appeared in numerous works in the sampling literature for example see Brosse et al. (2019), Lovas et al. (2023), Lytras and Sabanis (2025), Neufeld et al. (2022),Johnston et al. (2024) Lytras and Mertikopoulos (2025).
The drift coefficient will be designed based on a construction borrowed from the numerics of SDEs literature ( Johnston and Sabanis (2024)) , and the preserved monotonicity and Lipschtiz assumption are a major technical contribution compared to these earlier works.
@@
  • $h_{\lambda}(x)=h(x)\quad\forall x\in B(0,r_{\lambda}-2)$
  • $|h_{\lambda}(x)|\leq\frac{2}{\sqrt{\lambda}}(1+|x|)\quad\forall x\in\mathbb{R}^{d}$
  • $|h_{\lambda}(x)-h_{\lambda}(y)|\leq\frac{1}{\sqrt{\lambda}}|x-y|\quad\forall x,y\in\mathbb{R}^{d}$
  • $\langle h_{\lambda}(x)-h_{\lambda}(y),x-y\rangle\geq m|x-y|^{2}\quad\forall x,y\in\mathbb{R}^{d}$
@@ 4 Main results
$W_{2}(\mathcal{L}(\theta_{n}),\pi)\leq W_{2}(\mathcal{L}(\theta_{n},V_{n}),\Pi)\leq 2\sqrt{M_{\lambda}}(1-f(\lambda))^{n/2}W_{2}(\mathcal{L}(\theta_{n},V_{n}),\Pi)+C\lambda^{2}/f(\lambda)$
@@
$a\ =\ 1/M_{\lambda}\,,\qquad b\ =\ 1/\gamma\,,\qquad\kappa\ =\ m/(3\gamma)\,.$
@@
Let $(x_{n},v_{n})$ be the $n$-th iterate of the OBABO tamed scheme with stepsize $\lambda$. Then, there holds
$W_{2}(\mathcal{L}(x_{n}),\pi)\leq W_{2}(\mathcal{L}(x_{n},v_{n}),\Pi)\leq 3\sqrt{M_{\lambda}}(1-\frac{3m}{\gamma}\lambda)^{n/2}W_{2}(\mathcal{L}(x_{n},v_{n}),\Pi)+C^{\prime}M_{\lambda}\frac{\lambda\gamma}{m}$
@@ 4.1 Comparison with related literature
The contribution of the current article is twofold. On the one hand, it provides contraction rates for underdamped numerical schemes beyond global Lipschitz continuity of the drift. This is quite novel, as the common way to study these schemes is to study the contraction properties of the continuous type dynamics, which may sometimes be very complicated. For example in Johnston et al. (2024) a very complicated proof was employed which relied on the use of the Moreau-Yosida transform to essentially first sample from the Moreau-Yosida measure and the compare to the original. On the other hand, this work provides the first convergence results for the Velvet-scheme operator in the super-linearly growing case while also considerably improves the rate of convergence for a tamed exponential euler scheme than the one obtained in Johnston et al. (2024). Given the the fact that in many applications the underdamped scheme has been performing better than the overdamped our algorithm can be employed as a viable alternative to other overdamped tamed Langevin algorithms (Brosse et al. (2019),Lytras and Mertikopoulos (2024),Lytras and Sabanis (2025),Lovas et al. (2023),Neufeld et al. (2022)). In addition, another novel contribution of our work is the proof of Log-Sobolev inequalities for the iterates of algorithm. There has been a lot of active research in this domain, as such inequalites are quite important to establish results in differential privacy.
@@ 5 Contraction rates for different schemes-Strategy
Theorem 7 (Horn and Zhang (2005), Theorem 1.6).
@@
Throughout our proof a prominent role plays the control of the quantity $h_{\lambda}(x)-h_{\lambda}(y)\quad\forall x,y\in\mathbb{R}^{d}$. A key observation is the fact that if $h_{\lambda}$ is $L$-Lipschitz and $m$ strongly monotone (with the constant $L$ possibly depending on the stepsize, which will be discussed in the next section) then one obtains $h_{\lambda}(x)-h_{\lambda}(y)=Q(x-y)$ for some matrix, where $mI_{d}\leq Q\leq LI_{d}$. It is therefore crucial to have construct a drift coefficient which is global Lipschitz smooth but also strongly monotone.
@@ 6.2 Proof of Contraction rates of the Monotonic polygonal Stochastic exponential scheme
Now we will denote $\bar{x}_{j}:=\left(\tilde{x}_{j}-x_{j}\right),\bar{v}_{j}=\left(\tilde{v}_{j}-v_{j}\right)$ and $\bar{z}_{j}=\left(\bar{x}_{j},\bar{v}_{j}\right)$. Let $\eta=e^{-\lambda\gamma}$.
@@
$\bar{x}_{n+1}=\bar{x}_{n}+\frac{1-\eta}{\gamma}\bar{v}_{n}-\frac{\gamma\lambda+\eta-1}{\gamma^{2}}Q\bar{x}_{n},\quad\bar{v}_{n+1}=\eta\bar{v}_{n}-\frac{1-\eta}{\gamma}Q\bar{x}_{n},$
@@
$-\left(\frac{2b(1-\eta)(-1+\eta+\gamma\lambda)}{\gamma^{3}}+\frac{(-1+\eta+\gamma\lambda)^{2}}{\gamma^{4}}\right)\operatorname{eig}(Q^{2})$ $\geq-\left(\lambda^{2}\gamma^{2}-(1-\eta)^{2}\right)\frac{\operatorname{eig}(Q^{2})}{\gamma^{4}}$
@@ 6.3 Proof of Contraction rates for Tamed OBABO scheme
$a\ =\ 1/M_{\lambda}\,,\qquad b\ =\ 1/\gamma\,,\qquad\kappa\ =\ m/(3\gamma)\,.$
@@
$\begin{pmatrix}\bar{x}^{\prime}\\ \bar{v}^{\prime}\end{pmatrix}$ $=\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}+\begin{pmatrix}0&\lambda\tilde{\eta}I_{d}\\ -\frac{\lambda\tilde{\eta}}{2}(Q+Q^{\prime})&(\tilde{\eta}^{2}-1)I_{d}\end{pmatrix}\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}-\frac{\lambda}{2}\begin{pmatrix}\lambda Q\bar{x}\\ \tilde{\eta}Q^{\prime}\Delta\bar{x}\end{pmatrix}=(I_{2d}+\lambda A)\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}+h$
@@
$A=\begin{pmatrix}0&I_{d}\\ -\frac{1}{2}(Q+Q^{\prime})&-\gamma I_{d}\end{pmatrix}\,,\qquad h=\begin{pmatrix}0&\lambda(\tilde{\eta}-1)I_{d}\\ -\frac{\lambda(\tilde{\eta}-1)}{2}(Q+Q^{\prime})&(\tilde{\eta}^{2}-1+\lambda\gamma)I_{d}\end{pmatrix}\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}-\frac{\lambda}{2}\begin{pmatrix}\lambda Q\bar{x}\\ \tilde{\eta}Q^{\prime}\Delta\bar{x}\end{pmatrix}\,$.
@@
$\|h\|_{a,b}^{2}$ $\leqslant\frac{3}{2}\left|\lambda(\tilde{\eta}-1)\bar{v}-\frac{\lambda^{2}}{2}Q\bar{x}\right|^{2}+\frac{3}{2}a\left|-\frac{\lambda(\tilde{\eta}-1)}{2}(Q+Q^{\prime})\bar{x}+(\tilde{\eta}^{2}-1+\lambda\gamma)\bar{v}-\frac{\lambda\tilde{\eta}}{2}Q^{\prime}\Delta\bar{x}\right|^{2}$
$\leqslant 3\lambda^{4}\gamma^{2}/4|\bar{v}|^{2}+3\lambda^{4}M_{\lambda}^{2}/4|\bar{x}|^{2}+3a\left(\lambda^{2}\gamma M_{\lambda}/2+\lambda^{3}M_{\lambda}^{2}/4\right)^{2}|\bar{x}|^{2}+3a\left(\lambda^{2}\gamma^{2}/2+\lambda^{2}M_{\lambda}/2\right)^{2}|\bar{v}|^{2}$
@@
$\|A\bar{z}\|_{a,b}^{2}\ \leqslant\ \frac{3}{2}|\bar{v}|^{2}+\frac{3}{2}a\left|-\frac{1}{2}(Q+Q^{\prime})\bar{x}-\gamma\bar{v}\right|^{2}\ \leqslant\ 3M_{\lambda}|\bar{x}|^{2}+3a\left(\frac{M_{\lambda}}{2}+\gamma^{2}\right)|\bar{v}|^{2}\ \leqslant\ \frac{27\gamma^{2}}{8}\|\bar{z}\|_{a,b}^{2}\,$,

### v2

@@ 1 Introduction
$\mathrm{d}L_{t}$ $=-\nabla u\left(L_{t}\right)\mathrm{d}t+\sqrt{\frac{2}{\beta}}\mathrm{\penalty 10000\ d}B_{t},\qquad t\geq 0$
@@
and exploiting the fact that this SDE admits the target measure $\exp(-\beta u)$ as its unique invariant measure. Inspired by this observation Langevin-based algorithms have become a popular choice for sampling from distributions in high dimensional spaces. There has been a lot of work in providing non-asymptotic results for Unadjusted Langevin algorithm (ULA) under various assumptions such as Lipschitz continuity of $\nabla u$ and convexity of $u$. Under the assumption of convexity and gradient Lispchitz continuity important results are obtained in dalalyan2017theoretical; durmus2017nonasymptotic; durmus2019high; hola; convex, while in the non-convex case, under convexity at infinity or dissipativity assumptions, one may consult berkeley majka2020nonasymptotic, erdogdu2022convergence for ULA while for the Stochastic Gradient variant (SGLD) important works are raginsky, nonconvex, zhang2023nonasymptotic.
More recently, starting with the work of vempala2019rapid important estimates have been obtained under the assumption that the target measure $\pi$ satisfies an isoperimetric inequality and the gradient of $u$ satisfies a global Lipschitz continuity, (mou2022improved, balasubramanian2022towards). The latter assumption has also been relaxed to a weakly smooth assumption (with the gradient still satisfying a linear growth property), see nguyen2021unadjusted and erdogdu2021convergence.
Recent advances on sampling has led to a connection with optimization (vempala2019rapid,chewi2024ballistic) where overdamped Langevin dynamics can be seen as an analogous to gradient descent. It is therefore natural to explore momentum-like algorithms (nesterov1983method) for possible improvements. An example of such a method is the discretization of the underdamped/kinetic Langevin SDE given as
@@
where $\gamma>0$ is the friction coefficient and $B_{t}$ is a $d$-dimensional Brownian motion. Similarly to the overdamped Langevin SDE, the discretization of this diffusion can be used as both an MCMC sampler and non-convex optimizer ( gao2021global) since under appropriate conditions, the Markov process $\left\{\left(\tilde{\theta}_{t},\tilde{V}_{t}\right)\right\}_{t\geq 0}$ has a unique invariant measure given by
@@
Consequently, the marginal distribution of (2) in $\theta$ is precisely the target measure $\pi$. This means that sampling from (2) in the extended space and then keeping the samples in the $\theta$ space defines a valid sampler for the density $\pi$. The underdamped Langevin MCMC has shown improved convergence rates in the case of convex and gradient Lipschitz smooth potential $u$, e.g see cheng2018underdamped, dalalyan2018sampling,monmarche2021highaltschuler2024faster. In the non-convex setting important work has been done in Deniz, Gao, chau2019stochastic,10.3150/20-BEJ1297,zhang2023improved, while there has been a lot of effort in analyzing the contractive behaviour of these algorithms (monmarche2021high, leimkuhler2024contraction).
@@ 2.1 Blanket assumptions
$\langle h(x)-h(y),x-y\rangle\geq 2m|x-y|^{2}\qquad x,y\in\mathbb{R}^{d}.$
@@
$\mathbb{E}|\theta_{0}|^{2k}+\mathbb{E}|V_{0}|^{2k}<\infty\qquad\forall k\in\mathbb{N}.$
@@ 2.2.1 Stochastic exponential scheme
In this section we propose a modified version of the stochastic exponential scheme, an algorithm which was first developed in cheng2018underdamped and analysed in depth in dalalyan2018sampling and gao2021global under the assumption of Lipschitz continuity for the gradient. The creation of this algorithm is motivated by the fact that (1), after applying Itô’s formula to the product $e^{\gamma t}V_{t}$ and by following standard calculations, can be rewritten as
@@
  • for any fixed $j$, the random vectors $\left(\boldsymbol{\Xi}_{j}\right)_{1},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{1},\left(\boldsymbol{\Xi}_{j}\right)_{2},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{2},\ldots,\left(\boldsymbol{\Xi}_{j}\right)_{d},\left(\boldsymbol{\Xi}_{j}^{\prime}\right)_{d}$ are iid with covariance matrix
@@
$\mathcal{L}(\bar{Y}^{\lambda}_{n},\bar{x}^{\lambda}_{n})=\mathcal{L}(\tilde{Y}_{n\lambda},\tilde{x}_{n\lambda}),\qquad\forall n\in\mathbb{N}.$
@@ 3 Monotonicity preserving taming scheme
Taming schemes have appeared in numerous works in the sampling literature for example see tula, TUSLA, lytras2023taming, neufeld2022non,johnston2023kinetic pmlr-v258-lytras25a.
The drift coefficient will be designed based on a construction borrowed from the numerics of SDEs literature ( johnston2024strongly) , and the preserved monotonicity and Lipschtiz assumption are a major technical contribution compared to these earlier works.
@@
  • $h_{\lambda}(x)=h(x)\qquad\forall x\in B(0,r_{\lambda}-2)$
  • $|h_{\lambda}(x)|\leq\frac{2}{\sqrt{\lambda}}(1+|x|)\qquad\forall x\in\mathbb{R}^{d}$
  • $|h_{\lambda}(x)-h_{\lambda}(y)|\leq\frac{1}{\sqrt{\lambda}}|x-y|\qquad\forall x,y\in\mathbb{R}^{d}$
  • $\langle h_{\lambda}(x)-h_{\lambda}(y),x-y\rangle\geq m|x-y|^{2}\qquad\forall x,y\in\mathbb{R}^{d}$
@@ 4 Main results
$W_{2}(\mathcal{L}(\theta_{n}),\pi)\leq W_{2}(\mathcal{L}(\theta_{n},V_{n}),\Pi)\leq 2\sqrt{M_{\lambda}}(1-f(\lambda))^{n/2}W_{2}(\mathcal{L}(\theta_{0},V_{0}),\Pi)+C\lambda^{2}/f(\lambda)$
@@
$a\ =\ 1/M_{\lambda}\,,\hskip 18.49988ptb\ =\ 1/\gamma\,,\hskip 18.49988pt\kappa\ =\ m/(3\gamma)\,.$
@@
Let $(x_{n},v_{n})$ be the $n$-th iterate of the OBABO tamed scheme with stepsize $\lambda$. Then, we have
$W_{2}(\mathcal{L}(x_{n}),\pi)$ $\leq W_{2}(\mathcal{L}(x_{n},v_{n}),\Pi)$ (10)
$\leq 3\sqrt{M_{\lambda}}(1-\frac{3m}{\gamma}\lambda)^{n/2}W_{2}(\mathcal{L}(x_{0},v_{0}),\Pi)+C^{\prime}M_{\lambda}\frac{\lambda\gamma}{m}$
@@ 4.1 Comparison with related literature
The contribution of the current article is twofold. On the one hand, it provides contraction rates for underdamped numerical schemes beyond global Lipschitz continuity of the drift. This is quite novel, as the common way to study these schemes is to study the contraction properties of the continuous type dynamics, which may sometimes be very complicated. For example in johnston2023kinetic a very complicated proof was employed which relied on the use of the Moreau-Yosida transform to essentially first sample from the Moreau-Yosida measure and the compare to the original. On the other hand, this work provides the first convergence results for the Velvet-scheme operator in the super-linearly growing case while also considerably improves the rate of convergence for a tamed exponential euler scheme than the one obtained in johnston2023kinetic. Given the the fact that in many applications the underdamped scheme has been performing better than the overdamped our algorithm can be employed as a viable alternative to other overdamped tamed Langevin algorithms (tula,lytras2024tamed,lytras2023taming,TUSLA,neufeld2022non). In addition, another novel contribution of our work is the proof of Log-Sobolev inequalities for the iterates of algorithm. There has been a lot of active research in this domain, as such inequalites are quite important to establish results in differential privacy.
@@ 5 Contraction rates for different schemes-Strategy
Theorem 7 (horn2005basic, Theorem 1.6).
@@
Throughout our proof a prominent role plays the control of the quantity $h_{\lambda}(x)-h_{\lambda}(y)\qquad\forall x,y\in\mathbb{R}^{d}$. A key observation is the fact that if $h_{\lambda}$ is $L$-Lipschitz and $m$ strongly monotone (with the constant $L$ possibly depending on the stepsize, which will be discussed in the next section) then one obtains $h_{\lambda}(x)-h_{\lambda}(y)=Q(x-y)$ for some matrix, where $mI_{d}\leq Q\leq LI_{d}$. It is therefore crucial to have construct a drift coefficient which is global Lipschitz smooth but also strongly monotone.
@@ 6.1 Proof of properties of taming scheme
Proof of 2.
Proof of (1).
Noticing that $t(x)=1$ and $s(x)=0\qquad\forall x\in B(0,r_{\lambda}-2),$ it follows that
$h_{\lambda}(x)=g_{\lambda}(x)+mx=g(x)+mx=h(x)\qquad\forall x\in B(0,r_{\lambda}-2).$
Proof of (3).
We define $A_{1}=\bar{B}(0,r_{\lambda}-2)$, $A_{2}=\bar{B}(0,r_{\lambda}-1)/{B}(0,r_{\lambda}-2)$, $A_{3}=\bar{B}(0,r_{\lambda})/B_{r_{\lambda}-1}$, and $A_{4}=B(0,r_{\lambda})^{c}$. Since $g_{\lambda}=g$, on $A_{1}$ we have:
$|g_{\lambda}(x)-g_{\lambda}(y)|\leq Lip_{B_{r}}|x-y|.$
For $x,y\in A_{2}$ $Lip(s(x))=1$ and $t(x)=1$ on $A_{2}$ one obtains
For $x,y\in A_{3}$ since $Lip(t)=1$ and
$|g(x)|\leq R_{\lambda}\qquad\forall x\in A_{3}$
then,
$\geq R_{\lambda}^{2}(|x|-|y|)^{2}$
$\geq 0$
so for each set $A_{i}$ there holds
$|g_{\lambda}(x)-g_{\lambda}(y)|$ $=|g_{\lambda}(l_{0})-g_{\lambda}(l_{k})|$
$\leq\sum_{i=1}^{k}|g(l_{i-1})-g(l_{i})|$
$\leq(Lip_{B_{r}}+R_{\lambda}r_{\lambda})sum_{i=1}^{k}|l_{i}-l_{i-1}|$
$=(Lip_{B_{r}}+R_{\lambda}r_{\lambda})|x-y|$
where the last line holds since the segments are collinear. ∎
Proof of (4).
$\geq R_{\lambda}(s(x)|x|-s(y)|y|)(|x|-|y|)$
$\geq 0.$
For $x,y\in A_{3},$
$\langle g_{\lambda}(x)-g_{\lambda}(y),x-y\rangle$ $=R_{\lambda}|x-y|^{2}+t(y)\langle g(x)-g(y),x-y\rangle$
$\hskip 18.49988pt+(t(x)-t(y))\rangle g(x),x-y\rangle$
[…]
$\geq(R-|g(x)|)|x-y|^{2}\geq 0$
For $x,y\in A_{4}$,
$\langle g_{\lambda}(x)-g_{\lambda}(y),x-y\rangle=\sum_{i=1}^{k}\langle g_{\lambda}(l_{i})-g_{\lambda}(l{i-1}),l_{i}-l_{i-1}\rangle\geq 0.\qed$
This proves our claim and completes our proof. ∎
@@ 6.2 Proof of Contraction rates of the Monotonic polygonal Stochastic exponential scheme
Let $\bar{x}_{j}:=\left(\tilde{x}_{j}-x_{j}\right)$, $\bar{v}_{j}=\left(\tilde{v}_{j}-v_{j}\right)$ and $\bar{z}_{j}=\left(\bar{x}_{j},\bar{v}_{j}\right)$. Let $\eta=e^{-\lambda\gamma}$.
@@
$\bar{x}_{n+1}=\bar{x}_{n}+\frac{1-\eta}{\gamma}\bar{v}_{n}-\frac{\gamma\lambda+\eta-1}{\gamma^{2}}Q\bar{x}_{n},\qquad\bar{v}_{n+1}=\eta\bar{v}_{n}-\frac{1-\eta}{\gamma}Q\bar{x}_{n},$
@@
$-\left(\frac{2b(1-\eta)(-1+\eta+\gamma\lambda)}{\gamma^{3}}+\frac{(\eta+\gamma\lambda-1)^{2}}{\gamma^{4}}\right)\operatorname{eig}(Q^{2})$ $\geq\left((1-\eta)^{2}-\lambda^{2}\gamma^{2}\right)\frac{\operatorname{eig}(Q^{2})}{\gamma^{4}}$
@@ 6.3 Proof of Contraction rates for Tamed OBABO scheme
$a\ =\ 1/M_{\lambda}\,,\hskip 18.49988ptb\ =\ 1/\gamma\,,\hskip 18.49988pt\kappa\ =\ m/(3\gamma)\,.$
@@
$\begin{pmatrix}\bar{x}^{\prime}\\ \bar{v}^{\prime}\end{pmatrix}=\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}$ $+\begin{pmatrix}0&\lambda\tilde{\eta}I_{d}\\ -\frac{\lambda\tilde{\eta}}{2}(Q+Q^{\prime})&(\tilde{\eta}^{2}-1)I_{d}\end{pmatrix}\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}$
$-\frac{\lambda}{2}\begin{pmatrix}\lambda Q\bar{x}\\ \tilde{\eta}Q^{\prime}\Delta\bar{x}\end{pmatrix}=(I_{2d}+\lambda A)\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}+h$
@@
$A=\begin{pmatrix}0&I_{d}\\ -\frac{1}{2}(Q+Q^{\prime})&-\gamma I_{d}\end{pmatrix}$
and
$h=\begin{pmatrix}0&\lambda(\tilde{\eta}-1)I_{d}\\ -\frac{\lambda(\tilde{\eta}-1)}{2}(Q+Q^{\prime})&(\tilde{\eta}^{2}-1+\lambda\gamma)I_{d}\end{pmatrix}\begin{pmatrix}\bar{x}\\ \bar{v}\end{pmatrix}-\frac{\lambda}{2}\begin{pmatrix}\lambda Q\bar{x}\\ \tilde{\eta}Q^{\prime}\Delta\bar{x}\end{pmatrix}\,$.
@@
$\|h\|_{a,b}^{2}$ $\leqslant\frac{3}{2}\left|\lambda(\tilde{\eta}-1)\bar{v}-\frac{\lambda^{2}}{2}Q\bar{x}\right|^{2}$
$\hskip 18.49988pt+\frac{3}{2}a\left|-\frac{\lambda(\tilde{\eta}-1)}{2}(Q+Q^{\prime})\bar{x}+(\tilde{\eta}^{2}-1+\lambda\gamma)\bar{v}-\frac{\lambda\tilde{\eta}}{2}Q^{\prime}\Delta\bar{x}\right|^{2}$
$\leqslant 3\lambda^{4}\gamma^{2}/4|\bar{v}|^{2}+3\lambda^{4}M_{\lambda}^{2}/4|\bar{x}|^{2}+3a\left(\lambda^{2}\gamma M_{\lambda}/2+\lambda^{3}M_{\lambda}^{2}/4\right)^{2}|\bar{x}|^{2}$
$\hskip 18.49988pt+3a\left(\lambda^{2}\gamma^{2}/2+\lambda^{2}M_{\lambda}/2\right)^{2}|\bar{v}|^{2}$
@@
$\|A\bar{z}\|_{a,b}^{2}$ $\leqslant\frac{3}{2}|\bar{v}|^{2}+\frac{3}{2}a\left|-\frac{1}{2}(Q+Q^{\prime})\bar{x}-\gamma\bar{v}\right|^{2}\$
$\leqslant 3M_{\lambda}|\bar{x}|^{2}+3a\left(\frac{M_{\lambda}}{2}+\gamma^{2}\right)|\bar{v}|^{2}$
$\leqslant\frac{27\gamma^{2}}{8}\|\bar{z}\|_{a,b}^{2}\,$,